# Fluttering without wind: Stokesian quasiperiodic settling

### v1

@@ Abstract
Fluid inertia imparts rich dynamics to leaves falling from trees. We show that bodies with two planes of symmetry can display a range of behaviours even without inertia. Any such body supports a conserved quantity in its dynamics, and is either a settler, a drifter or a flutterer, depending only on its shape. At large time, settlers and drifters respectively fall vertically and obliquely, while flutterers rotate forever while executing intricate patterns. The dynamics of flutterers decouples into a periodic and a Floquet part with different time scales, giving periodicity or quasiperiodicity. We design a set of bodies and use the boundary integral method to show that settlers, drifters and flutterers, all lie in this set.
@@
The complex and poetic dance of falling autumn leaves is a high Reynolds number phenomenon, which depends crucially on vortex shedding, a feature absent in Stokes flow [1, 2, 3]. One would therefore expect settling at zero Reynolds number to be uneventful, especially if the body has a high degree of symmetry in its shape. In fact, in steady Stokes flow, the linear and angular motion of a body with three planes of symmetry, like a sphere or an ellipsoid, are completely decoupled, and in absence of an external torque it can only display linear motion at constant velocity [4, 5, 6]. Bodies with fewer symmetries, however, can couple their translation and rotation dynamics, opening up richer possibilities. One of the consequences of this coupling is the chiral sedimenting trajectories observed for chiral body shapes, such as helices and Kelvin’s isotropic helicoid. [7, 8, 9].
Apart from complexity in body shape, another way of coupling rotation and translation in a body is through hydrodynamic interactions with other bodies [4, 5, 10, 11]. For example, two spheroids have been shown [12] to display periodic sedimentation in direct analogy with Kepler’s orbit.
A coupling between translation and rotation is not the only way to obtain non-trivial Stokesian dynamics, even in a single body. An externally imposed background flow such as shear can achieve this. The angular velocity provided by simple shear can cause an ellipsoid, despite its high symmetry, and also a body with two planes of symmetry, to go into chaotic orbits [13, 14].
@@
Our interest is in the sedimentation of a single achiral body in zero background flow. Ellipsoids have been well-studied, and we go a step further by letting go of only one of the symmetries. We opt for bodies with two planes of symmetry, which support a particularly simple translation-rotation coupling, unlike ellipsoids, which have none. We show that all bodies of this description fall into one of three classes: settlers, drifters and flutterers, as we term them. Settlers asymptotically adopt an orientation and fall vertically, while drifters asymptotically attain a constant velocity with both horizontal and vertical components. Flutterers are the most interesting, showing quasiperiodic or periodic, but never chaotic, motion. We show that the dynamics is completely described by an overlaying of Floquet dynamics in one of the horizontal projections of the body-fixed coordinates on the periodic dynamics of the vertical projections. Our predictions are supported by recent experiments on a U-shaped disk [15] which falls into our class of flutterers. We design a set of bodies containing settlers, drifters and flutterers, and obtain their resistance matrix, which relates the external forces to particle motion, numerically by the boundary integral method. The design involves replacing the circular cross-section of a bent capsule by a concavo-convex one (see figure 1). We shall return to this specific set of bodies later but our derivation and findings hold for any body with two planes of symmetry.
@@
In the over-damped limit where we work, gravity acting on the body is balanced by the viscous drag. The Stokesian dynamics is conveniently written in the body-fixed coordinate system of unit vectors $\{\bm{p}_{1},\bm{p}_{2},\bm{p}_{3}\}$ shown in figure 1, as
$\begin{pmatrix}F_{1}\\ F_{2}\\ F_{3}\\ 0\\ 0\\ 0\\ \end{pmatrix}=\underbrace{\begin{pmatrix}A_{1}&0&0&0&0&B_{31}\\ 0&A_{2}&0&0&0&0\\ 0&0&A_{3}&B_{13}&0&0\\ 0&0&B_{13}&C_{1}&0&0\\ 0&0&0&0&C_{2}&0\\ B_{31}&0&0&0&0&C_{3}\\ \end{pmatrix}}_{\cal R}\begin{pmatrix}U_{1}\\ U_{2}\\ U_{3}\\ \Omega_{1}\\ \Omega_{2}\\ \Omega_{3}\\ \end{pmatrix},$ (1)
@@
where $\bm{F}=F_{1}\bm{p}_{1}+F_{2}\bm{p}_{2}+F_{3}\bm{p}_{3}$ is the external force, and $\bm{U}=U_{1}\bm{p}_{1}+U_{2}\bm{p}_{2}+U_{3}\bm{p}_{3}$ and $\bm{\Omega}=\Omega_{1}\bm{p}_{1}+\Omega_{2}\bm{p}_{2}+\Omega_{3}\bm{p}_{3}$ the translational and angular velocities respectively. Note that the external torque about the center of mass is zero. Equation 1 has been is non-dimensionalized by characteristic length and time scales $L$ and $\tau$ of the body, with $\tau=\eta L^{2}/(mg_{r})$, $\eta$ being the fluid viscosity, and $mg_{r}$ the buoyancy-corrected weight of the body. All entries in the resistance matrix $\cal R$ depend only on the body shape and are invariant under overall size-scaling due to our non-dimensionalisation. We note that the two planes of symmetry allow for translation-rotation coupling of a particularly simple form, with two independent entries $B_{13}$ and $B_{31}$. The rotation matrix $\bm{R}$ connects the lab-fixed coordinate axes to the body-fixed coordinate axes as: $\bm{p}_{i}=\bm{R}_{ji}\bm{e}_{j}$, $i,j\in\{1,2,3\}$ with $\bm{e}_{1}=\bm{\hat{x}},\bm{e}_{2}=\bm{\hat{y}}$ and $\bm{e}_{3}=\bm{\hat{z}}$. So, in the lab frame
@@
Figure 1: (a) A saddle-shaped body with two mutually perpendicular planes of symmetry, with principal axes $\bm{p}_{1}$, $\bm{p}_{2}$ and $\bm{p}_{3}$. The lab-fixed coordinate axes are $x$, $y$ and $z$ axes, $y$ being anti-parallel to gravity $\bm{g}$. (b) Note that $\bm{p}_{1}-\bm{p}_{3}$ is not a plane of symmetry.
We define two parameters $\alpha_{p}$ and $\alpha_{r}$ as:
@@
which are functions only of the shape of the body. The pitching rate $\Omega_{1}$ and the rolling rate $\Omega_{3}$ are proportional to the two parameters $\alpha_{p}$ and $\alpha_{r}$, respectively. We obtain a fortunate decoupling of the dynamics of the vertical projections of the body’s coordinates from that of their horizontal projections. However, the horizontal projections depend on the vertical projections, as we shall see. For the vertical projections, using equations (1), (5) and $\bm{\dot{p}}_{i}=\bm{\Omega}\times\bm{p}_{i},\,i\in\{1,2,3\}$, we get
@@
with the constraint $|\bm{p}_{y}|^{2}=1$, where $\bm{p}_{y}\equiv(p_{1y},p_{2y},p_{3y})$, as required by equation (5). Equation (Fluttering without wind: Stokesian quasiperiodic settling) admits a conserved quantity given by
@@
The great circles given by $p_{1y}=0$ and $p_{3y}=0$ are invariant solutions of equation (Fluttering without wind: Stokesian quasiperiodic settling), which means that the sign of $p_{1y}$ and $p_{3y}$ is preserved throughout the dynamics of $\bm{p}_{y}$. We have exploited this to define $H$ in terms of absolute values of $p_{1y}$ and $p_{3y}$ to ensure that $H$ remains a positive real number. The conserved quantities $H$ and $|\bm{p}_{y}|^{2}=1$ render the system (Fluttering without wind: Stokesian quasiperiodic settling) integrable. The solutions lie on the intersection of the unit sphere $\mathcal{S}^{2}$ with surfaces of constant $H$.
The dynamical system described by equation (Fluttering without wind: Stokesian quasiperiodic settling) has fixed points given by:
@@
Figure 2: Phase portraits of the $\bm{p}_{y}$ dynamical system. Red, green and black dots denote unstable, stable and center fixed points, respectively, and blue lines with arrows are sample trajectories. (a) Settlers eventually fall vertically down with their $\bm{p}_{2}$ axis parallel/anti-parallel to gravity. (b) Drifters at long times sediment with horizontal drift, with their principal axes oriented obliquely with gravity. The beige plane represents a particular $H$. (c) Flutterers rotate forever as they fall, performing periodic or quasi-periodic orbits.
@@
Bodies with $\alpha_{p}\alpha_{r}<0$ are termed ‘settlers’, whose phase portrait has one stable and one unstable fixed point. From any initial orientation, settlers attain their stable orientation, with their $\bm{p}_{2}$ parallel/anti-parallel to gravity, asymptotically in time, and thence fall vertically. The stable fixed point is globally attracting (see appendix A). A bent capsule with circular cross-section, for example, is a settler.
@@
The class we term ‘drifters’ have $\alpha_{p}\alpha_{r}=0$. Their dynamics sustains infinitely many fixed points lying along the great circle, which we term $\cal A$, of $p_{1y}=0$ if $\alpha_{p}=0$ or $p_{3y}=0$ if $\alpha_{r}=0$. The dynamics occur in the intersection of $\mathcal{S}^{2}$ with the plane of constant $p_{3y}$ ($p_{1y}$), provided $\alpha_{p}=0$ ($\alpha_{r}=0$). This invariant manifold (a circle) contains two points of intersection with $\cal A$, one stable and the other unstable, as shown in figure 2(b). Drifters will eventually fall with their principal axes inclined at some constant angles with respect to gravity. Thus, like ellipsoids, drifters display persistent horizontal drift.
@@
A wide range of behaviour, all involving rotation for all time, is displayed by ‘flutterers’, whose $\alpha_{p}\alpha_{r}>0$. Every fixed point here is either a center or a saddle point. The maximum in $H$,
$H_{\text{max}}=\dfrac{|\alpha_{p}|^{|\alpha_{p}|/2}|\alpha_{r}|^{|\alpha_{r}|/2}}{(|\alpha_{p}|+|\alpha_{r}|)^{|\alpha_{p}|/2+|\alpha_{r}|/2}},$ (10)
@@
corresponds to center fixed points and the minimum, $H_{\text{min}}=0$, corresponds to stable and unstable manifolds of the saddle points. The trajectory for a given initial condition lies on one of the four closed curves formed by intersection of $\mathcal{S}^{2}$ with the corresponding constant $H$ surface (figure 3),
@@
Figure 3: (a) Contour plot of the conserved quantity $H$. Its maximum, $H_{\text{max}}$, corresponds to fixed points of the $\bm{p}_{y}$ dynamics (black dots). The great circles in red, where $H=0$, are the stable and unstable manifolds of the saddle points denoted by red dots, which divide $\mathcal{S}^{2}$ into four regions. Each trajectory of constant $H$ lives on one of the four closed regions. (b) Trajectories on the yellow surfaces with $p_{1y}p_{3y}>0$ have different translational chirality from those with $p_{1y}p_{3y}<0$ which lie on the blue surfaces, as discussed later.
@@
indicating periodic behaviour of $\bm{p}_{y}$. The time period $T_{y}$ of orbits near the center fixed points can be obtained after linearizing equation (Fluttering without wind: Stokesian quasiperiodic settling) about them: $\lim_{H\to H_{\text{max}}}T_{y}\equiv T_{ym}=2\pi/\sqrt{2\alpha_{p}\alpha_{r}}$. Also $T_{y}$ everywhere can be obtained numerically, and is shown in figure 4. Flutterers of different shapes show qualitatively similar trends in time period.
@@
Figure 4: Dependence of the time period of the $\bm{p}_{y}$ dynamics of flutterers on the conserved quantity $H$. As $H\to H_{\text{max}}$, $T_{y}\to T_{ym}$. As $H\to 0$, trajectories spend longer durations close to the saddle points, and $T_{y}$ can increase without bound. A given ratio $\alpha_{p}/\alpha_{r}$ and its inverse yield the same curve.
The equation (Fluttering without wind: Stokesian quasiperiodic settling) only describes the rotational dynamics up to a rotation about the $y$-axis. To capture the remaining part of the dynamics we write equations for the projection of $\bm{p}_{i}$’s along the $x$-axis of the lab frame:
$\begin{pmatrix}\dot{p}_{1x}\\ \dot{p}_{2x}\\ \dot{p}_{3x}\end{pmatrix}=\bm{A}(t)\begin{pmatrix}p_{1x}\\ p_{2x}\\ p_{3x}\end{pmatrix},$ (11g)
$\bm{A}(t)\equiv\begin{pmatrix}0&&\alpha_{r}p_{1y}(t)&&0\\ -\alpha_{r}p_{1y}(t)&&0&&\alpha_{p}p_{3y}(t)\\ 0&&-\alpha_{p}p_{3y}(t)&&0\\ \end{pmatrix}.$ (11k)
@@
$\bm{p}_{x}(t)\cdot\bm{p}_{y}(t)=0,\quad\forall\quad t.$ (12c)
Equations (Fluttering without wind: Stokesian quasiperiodic settling), (Fluttering without wind: Stokesian quasiperiodic settling) and (Fluttering without wind: Stokesian quasiperiodic settling) fully describe the rotational dynamics of the body. Equation (Fluttering without wind: Stokesian quasiperiodic settling) shows that $\bm{p}_{x}$ is driven by $\bm{p}_{y}$. Since the latter is time-periodic, $\bm{p}_{x}(t)$ can be solved using Floquet theory [16]. We define a $T_{y}$ period Poincare map of the $\bm{p}_{x}$ dynamical system as:
@@
where $\bm{\Phi}_{t_{0}}^{t}$ is the solution operator of equation (Fluttering without wind: Stokesian quasiperiodic settling), i.e., $\bm{p}_{x}(t)=\bm{\Phi}_{t_{0}}^{t}\bm{p}_{x}(t_{0})$. Since $\bm{A}^{T}(t)=-\bm{A}(t)$, equation (Fluttering without wind: Stokesian quasiperiodic settling) ensures that $\bm{\Phi}_{t_{0}}^{t}\in SO(3)$. Therefore, both the solution operator and the Poincare map correspond to some 3D rotations. Consequently, the eigenvalues of the Poincare map are given by $1$ and $e^{\pm i\mu T_{y}}$, where $\mu\in\mathbb{R}$ is the Floquet exponent. The constraints of (Fluttering without wind: Stokesian quasiperiodic settling) restrict $\bm{p}_{x}(t)$ to lie normal to $\bm{p}_{y}(t)$. Since $\bm{p}_{y}(t_{0}+T_{y})=\bm{p}_{y}(t_{0})$, the image set of the Poincare map $\{\mathcal{P}_{t_{0}}^{t_{0}+nT_{y}}\bm{p}_{x}(t_{0})\}_{n\in\mathbb{N}}$ is a subset of the great circle lying normal to $\bm{p}_{y}(t_{0})$. Thus, the Poincare map $\mathcal{P}_{t_{0}}^{t_{0}+T_{y}}$ acting on $\bm{p}_{x}$ corresponds to the rotation of $\bm{p}_{x}$ by an angle $\theta\equiv\mu T_{y}$ about the axis $\bm{p}_{y}(t_{0})$. It is shown in the supplementary material that the eigenvalues of the Poincare map are independent of the initial time $t_{0}$. So the rotation angle $\theta=\mu T_{y}$ is only a function of $H$. By Floquet theory, any solution of equation (Fluttering without wind: Stokesian quasiperiodic settling) can be decomposed as
@@
Figure 5: The $\bm{p}_{x}$ dynamics of flutterers involves two frequencies, $\mu$ and $\omega_{y}$. Their ratio, which is equivalent to the rotation angle $\theta/2\pi$ of the Poincare map, is shown here.
@@
There are differences in detail for different ratios of $\alpha_{r}$ and $\alpha_{p}$, but in all cases, the frequency ratio varies continuously on the interval [11/2, $[1-1/\sqrt{2},\,0.5]$ which is therefore dense in periodic and quasi-periodic orbits. Figure 6 shows phase trajectories and image sets of Poincare maps for a periodic and quasi-periodic case.
@@
Figure 6: Sample $\bm{p}_{x}$ trajectories shown as black curves on $\mathcal{S}^{2}$. (a) Period-5 trajectory. The red dots show the Poincare map $\mathcal{P}_{t_{0}}^{t_{0}+T_{y}}$ which lie on the great circle shown by the blue curve. This great circle lies normal to $\bm{p}_{y}(t_{0})$. (b) Quasi-periodic trajectory. The Poincare map covers the entire great circle lying normal to $\bm{p}_{y}(t_{0})$. The filled region has the inversion symmetry $\bm{p}_{x}\to-\bm{p}_{x}$.

### v2

# Sedimentation Dynamics of Bodies with Two Planes of Symmetry

@@ Abstract
We show that bodies with two planes of symmetry can display a range of behaviours even without inertia. Any such body supports a conserved quantity in its dynamics, and is either a settler, a drifter or a flutterer, depending only on its shape. At large time, settlers and drifters respectively fall vertically and obliquely, while flutterers rotate forever while executing intricate patterns. The dynamics of flutterers decouples into a periodic and a Floquet part with different time scales, giving periodicity or quasi-periodicity. We design a set of bodies and use the boundary integral method to show that settlers, drifters and flutterers, all lie in this set.
DOI: 10.1103/PhysRevLett.134.014002
@@
The complex and poetic dance of falling autumn leaves is a high Reynolds number phenomenon, which depends crucially on vortex shedding [1, 2, 3], a feature absent in Stokes flow. One would therefore expect settling at zero Reynolds number to be uneventful, especially for a body shape of high symmetry. In fact, in steady Stokes flow, the linear and angular motion of bodies with three planes of symmetry, like ellipsoids, are completely decoupled, and without external torque can only display linear motion [4, 5, 6, 7, 8]. But particles can have complex shapes with fewer symmetries, and understanding their sedimentation is crucial in diverse contexts, from marine snow sequestering atmospheric carbon to ice crystals in clouds and industrial powder production, like milk powder. Unlike ellipsoids, bodies with fewer symmetries can couple translation and rotation, opening up richer possibilities, e.g., chiral sedimenting trajectories of chiral objects like helices and Kelvin’s isotropic helicoid [9, 10, 11]. The pioneering work of Brenner [12] examined the different possible terminal sedimentation behaviours allowed for various body symmetries. However, there are bodies which do not attain a steady terminal state. Recent work on sedimentation of a U-shaped disk [13] shows that achiral bodies can display quasi-periodic dynamics with chiral trajectories, and provides theory for the pitch and roll dynamics. This study also argues that the most complicated motion an arbitrary body can perform is quasi-periodic, and not chaotic. We present a general theory for the sedimentation of any body with two planes of symmetry, henceforth called di-bilaterals, and find three kinds of dynamics, the most interesting being quasi-periodic. The body shape allows for apriori classification in sedimentation behaviour. Our choice of dynamical variables brings out a new conserved quantity which restricts the dynamics of di-bilaterals. We explain the quasi-periodicity using Floquet theory.
Non-spherical particles are often modelled as ellipsoids to account for their orientation degrees of freedom. But ellipsoids sediment with persistent horizontal drift. We show that only a very narrow class of di-bilaterals, which includes ellipsoids, can show persistent drift. Moreover we prove that generic bodies displaying quasi-periodic motion cannot show persistent drift (see Supplemental Material). Our study is a strong indication that di-bilaterals, which can exhibit settling, drifting, helical and quasi-periodic motion, are a far better model of achiral bodies than ellipsoids, capable of explaining a range of sedimentation behaviour. Elastic fibers provide a natural setting for observing di-bilateral sedimentation [14].
@@
Our interest is in the sedimentation of a single achiral body in zero background flow. We opt for di-bilaterals, which support simple translation-rotation coupling (TRC), unlike ellipsoids. We show that all such bodies fall into one of three classes: settlers, drifters and flutterers. Settlers asymptotically align one of their principal axes along gravity and fall vertically, while drifters fall obliquely at constant speed. Flutterers are the most interesting, showing quasi-periodic or periodic, but never chaotic, motion. We show that their motion is completely described by an overlaying of Floquet dynamics in the horizontal, onto periodic dynamics in the vertical direction. We design a set of di-bilaterals (see figure 1) and obtain their resistance matrices, which relate the external forces to particle motion, numerically by the boundary integral method.
@@
In the over-damped limit where we work, gravity is balanced by viscous drag. The Stokesian dynamics is conveniently written in the body-fixed coordinate system of unit vectors $\{\bm{p}_{1},\bm{p}_{2},\bm{p}_{3}\}$ (see figure 1), as [4, 15, 13]:
$\begin{pmatrix}F_{1}\\ F_{2}\\ F_{3}\\ 0\\ 0\\ 0\\ \end{pmatrix}=\underbrace{\begin{pmatrix}A_{1}&0&0&0&0&B_{31}\\ 0&A_{2}&0&0&0&0\\ 0&0&A_{3}&B_{13}&0&0\\ 0&0&B_{13}&C_{1}&0&0\\ 0&0&0&0&C_{2}&0\\ B_{31}&0&0&0&0&C_{3}\\ \end{pmatrix}}_{\cal R}\begin{pmatrix}v_{1}\\ v_{2}\\ v_{3}\\ \Omega_{1}\\ \Omega_{2}\\ \Omega_{3}\\ \end{pmatrix},$ (1)
@@
where $\bm{F}=F_{1}\bm{p}_{1}+F_{2}\bm{p}_{2}+F_{3}\bm{p}_{3}$ is the external force, and $\bm{V}=v_{1}\bm{p}_{1}+v_{2}\bm{p}_{2}+v_{3}\bm{p}_{3}$ and $\bm{\Omega}=\Omega_{1}\bm{p}_{1}+\Omega_{2}\bm{p}_{2}+\Omega_{3}\bm{p}_{3}$ the translational and angular velocities respectively. Equation 1 has been non-dimensionalized by characteristic length and time scales $L$ and $\tau=\eta L^{2}/(mg_{r})$, $\eta$ being the fluid viscosity, and $mg_{r}$ the body’s buoyancy-corrected weight. The non-dimensional resistance matrix $\cal R$ depends only on the body shape. The geometry allows for TRC of a particularly simple form, with two independent entries $B_{13}$ and $B_{31}$. The lab-fixed and body-fixed coordinate axes are related by $\bm{p}_{i}=\bm{R}_{ji}\bm{e}_{j}$, $i,j\in\{1,2,3\}$ with $\bm{e}_{1}=\bm{\hat{x}},\bm{e}_{2}=\bm{\hat{y}}$ anti-parallel to gravity, and $\bm{e}_{3}=\bm{\hat{z}}$. So, in the lab frame
@@
Figure 1: (a) A body with two mutually perpendicular planes of symmetry, and the coordinate systems. (b) Showing that $\bm{p}_{1}-\bm{p}_{3}$ is not a plane of symmetry.
We define parameters
@@
which are functions only of the body shape. Rotational invariance in the body frame about the gravity axis provides, for any arbitrary-shaped body, a decoupling of the dynamics of the vertical projections of the body’s coordinate-axes from that of their horizontal projections [13, 16, 17, 6]. For the vertical projections, using equations (1), (2) and $\bm{\dot{p}}_{i}=\bm{\Omega}\times\bm{p}_{i},\,i\in\{1,2,3\}$, we get
@@
with the constraint $|\bm{p}_{y}|^{2}=1$, where $\bm{p}_{y}\equiv(p_{1y},p_{2y},p_{3y})$, as required by equation (2). Equation (4) admits a conserved quantity
@@
The relationship between $H$ and the continuous symmetries of (4) is discussed in the Supplemental Material [18]. Since $p_{1y}=0$ and $p_{3y}=0$ are invariant solutions of (4), the signs of $p_{1y}$ and $p_{3y}$ are preserved. We have exploited this to define $H$ to be a positive real number. The conserved quantities $H$ and $|\bm{p}_{y}|^{2}=1$ render the system (4), integrable, and restrict the solutions to go towards a fixed point or be periodic. Solutions lie on the intersection of the unit sphere $\mathcal{S}^{2}$ with surfaces of constant $H$. $H$ is a generalisation to arbitrary di-bilaterals, of a conserved quantity shown by [16, 6] for any body whose mobility centre coincides with the centre of mass, which for di-bilaterals is those with $\alpha_{p}=\alpha_{r}$.
Earlier studies [16, 17, 6] on the possible sedimentation dynamics of bodies of arbitrary shapes focused on the part of the dynamics described by the gravity vector viewed in the body frame. This dynamical quantity is equivalent to our $\bm{p}_{y}$. They showed that the $\bm{p}_{y}$ dynamics of a generic body can have stable, unstable, saddle and centre fixed points along with limit cycles. Our conserved quantity $H$, removes the possibility of a limit cycle in the $\bm{p}_{y}$ dynamics of di-bilaterals, but retains all other possible fixed points, as shown in figure 2.
Equation (4) supports the fixed points
@@
Figure 2: Phase portraits of the $\bm{p}_{y}$ dynamical system. Red, green and black dots denote unstable, stable and centre fixed points, respectively, and blue lines with arrows are sample trajectories. (a) Settlers and (b) drifters ultimately align their $\bm{p}_{2}$ axis along and obliquely to gravity respectively. The beige plane in (b) represents a particular $H$. (c) Flutterers rotate forever.
@@
Di-bilaterals with $\alpha_{p}\alpha_{r}<0$ are termed ‘settlers’, whose phase-portrait has one globally attracting (see Supplemental Material) stable fixed point and one unstable fixed point. Asymptotically in time, settlers orient their $\bm{p}_{2}$ parallel/anti-parallel to gravity and thence fall vertically, without rotating. Similar behaviour occurs in bent achiral fibres [19] and non-deformable chains [20]. Incidentally, a body lying in this class shows settler-like dynamics at moderate Reynolds number [21].
@@
The class we term ‘drifters’ have $\alpha_{p}\alpha_{r}=0$. Their dynamics supports infinitely many fixed points lying along the great circles $\cal A$, of $p_{1y}=0$ if $\alpha_{p}=0$ or $p_{3y}=0$ if $\alpha_{r}=0$ (figure 2(b)). The dynamics occur in the intersection of $\mathcal{S}^{2}$ with the plane of constant $p_{3y}$ ($p_{1y}$), provided $\alpha_{p}=0$ ($\alpha_{r}=0$). This invariant manifold (a circle) intersects $\cal A$, at a stable and an unstable fixed-point. Drifters eventually fall with their principal axes inclined at some constant angle to gravity and, like ellipsoids, display persistent horizontal drift and no rotation.
@@
A range of behaviour, all involving perpetual rotation, is displayed by ‘flutterers’, whose $\alpha_{p}\alpha_{r}>0$. Every fixed point here is either a centre or a saddle point, as for a U-shaped disk [22]. The maximum in $H$,
$H_{\text{max}}=\dfrac{|\alpha_{p}|^{\alpha_{p}/2}|\alpha_{r}|^{\alpha_{r}/2}}{|\alpha_{p}+\alpha_{r}|^{\alpha_{p}/2+\alpha_{r}/2}},$ (7)
@@
corresponds to centre fixed points and the minimum, $H_{\text{min}}=0$, corresponds to stable and unstable manifolds of the saddle points. The trajectory for a given initial condition lies on a closed curve representing a constant $H$ surface (figure 3),
@@
Figure 3: (a) Contour plot of the conserved quantity $H$ for flutterers. Closed curves are of constant $H$, with $H_{\text{max}}$ shown by the black dots, and $H=0$ corresponding to the red heteroclinic orbits of the saddle points (red dots). (b) Trajectories on the yellow surfaces with $p_{1y}p_{3y}>0$ have different translational chirality from those with $p_{1y}p_{3y}<0$ on the blue surfaces, as shown in the Supplemental Material.
@@
indicating periodic behaviour of $\bm{p}_{y}$. A similar periodic behavior in the inclination angles of a U-shaped disk relative to gravity has also been explained in the context of its sedimentation [13]. The time period $T_{y}$ of orbits near the centre fixed points can be obtained upon linearizing equation (4) about them: $\lim_{H\to H_{\text{max}}}T_{y}\equiv T_{ym}=2\pi/\sqrt{2\alpha_{p}\alpha_{r}}$. Also $T_{y}$ everywhere can be obtained numerically, and is shown in figure 4. Flutterers of different shapes show qualitatively similar trends in time period. Generically, displacing the mobility centre from the centre of mass leads to bifurcation in $\bm{p}_{y}$ dynamics [6, 17, 23]. However, such displacement in our flutterers preserves the nature of fixed points.
@@
Figure 4: Time period of the $\bm{p}_{y}$ dynamics of flutterers for different $H$. As $H\to 0$, $T_{y}$ can increase without bound. A given ratio $\alpha_{p}/\alpha_{r}$ and its inverse yield the same curve.
For the remaining part of the dynamics we write equations for the projection of $\bm{p}_{i}$’s along the $x$-axis:
$\begin{pmatrix}\dot{p}_{1x}\\ \dot{p}_{2x}\\ \dot{p}_{3x}\end{pmatrix}=\bm{A}(t)\begin{pmatrix}p_{1x}\\ p_{2x}\\ p_{3x}\end{pmatrix},$ (8a)
$\bm{A}(t)\equiv\begin{pmatrix}0&&\alpha_{r}p_{1y}(t)&&0\\ -\alpha_{r}p_{1y}(t)&&0&&\alpha_{p}p_{3y}(t)\\ 0&&-\alpha_{p}p_{3y}(t)&&0\\ \end{pmatrix}.$ (8b)
@@
$\bm{p}_{x}(t)\bm{\cdot}\bm{p}_{y}(t)=0,\quad\forall\quad t.$ (9c)
Equations (4), (8) and (9) fully describe the rotational dynamics of di-bilaterals. Equation (8) shows that $\bm{p}_{x}$ is driven by $\bm{p}_{y}$. Since the latter is time-periodic for flutterers, $\bm{p}_{x}(t)$ can be solved using Floquet theory [24]. We define a $T_{y}$ period Poincare map of the $\bm{p}_{x}$ dynamical system as:
@@
where $\bm{\Phi}_{t_{0}}^{t}$ is the solution operator of equation (8), i.e., $\bm{p}_{x}(t)=\bm{\Phi}_{t_{0}}^{t}\bm{p}_{x}(t_{0})$. Since $\bm{A}^{T}(t)=-\bm{A}(t)$, equation (10) ensures that $\bm{\Phi}_{t_{0}}^{t}\in SO(3)$. Therefore, both the solution operator and the Poincare map correspond to some 3D rotations. Consequently, the eigenvalues of the Poincare map are given by $1$ and $e^{\pm i\mu T_{y}}$, where $\mu\in\mathbb{R}$ is the Floquet exponent. The constraints of (9) restrict $\bm{p}_{x}(t)$ to lie normal to $\bm{p}_{y}(t)$. Since $\bm{p}_{y}(t_{0}+T_{y})=\bm{p}_{y}(t_{0})$, the image set of the Poincare map $\{\mathcal{P}_{t_{0}}^{t_{0}+nT_{y}}\bm{p}_{x}(t_{0})\}_{n\in\mathbb{N}}$ is a subset of the great circle lying normal to $\bm{p}_{y}(t_{0})$. Thus, the Poincare map $\mathcal{P}_{t_{0}}^{t_{0}+T_{y}}$ acting on $\bm{p}_{x}$ corresponds to the rotation of $\bm{p}_{x}$ by an angle $\theta\equiv\mu T_{y}$ about the axis $\bm{p}_{y}(t_{0})$. The Supplemental Material shows that $\theta$ depends only on $H$ and not on $t_{0}$. By Floquet theory, any solution of equation (8) can be decomposed as
@@
Figure 5: The $\bm{p}_{x}$ dynamics of flutterers involves two frequencies, $\mu$ and $\omega_{y}$. Their ratio, shown here, is the rotation angle $\theta/2\pi$ of the Poincare map. The black horizontal line is the lower limit $1-1/\sqrt{2}$ of the ratio.
@@
There are differences in detail for different flutterer shapes, but in all cases, the frequency ratio varies continuously on [11/2, $[1-1/\sqrt{2},\,0.5]$, showing dense existence of periodic and quasi-periodic orbits. Figure 6 shows phase trajectories and image sets of Poincare maps for a periodic and quasi-periodic case.
@@
Figure 6: Sample $\bm{p}_{x}$ trajectories shown as black curves on $\mathcal{S}^{2}$. (a) Period-5 trajectory. The red dots show the Poincare map $\mathcal{P}_{t_{0}}^{t_{0}+T_{y}}$ which lie on the great circle shown by the blue curve. This great circle lies normal to $\bm{p}_{y}(t_{0})$. (b) Quasi-periodic trajectory. The Poincare map covers the entire great circle lying normal to $\bm{p}_{y}(t_{0})$. The filled black region has the inversion symmetry $\bm{p}_{x}\to-\bm{p}_{x}$.